Discussions
bharat

· started a discussion

· 1 Months ago

probability at least 1 passes = 1 - Probability both fails
= 1 - 2/3*3/4
= 1- 1/2 = 1/2

Question:

The probability of A passing an exam is \(\cfrac{1}{3}\) and that of B passing the exam is \(\cfrac{1}{4}\). What is the probability that at least one of them passes the exam?

Options:
A) \(\cfrac{1}{12}\)
B) \(\cfrac{7}{12}\)
C) \(\cfrac{1}{2}\)
D) \(\cfrac{1}{3}\)
E) \(\cfrac{5}{12}\)
Solution:
Ans: (c) 

probability that at least one of them passes the exam = 1-(None is passed) = 1 - {(2/3)(3/4)} = 1-(1/2) = 1/2

Knowledge Expert

· commented

· 1 Months ago

option(c) is correct...

All Rights Reserved Top Rankers